OpenStudy (anonymous):

Is the function g(x)=x^3-6x odd, even, or neither?

OpenStudy (anonymous):

odd

OpenStudy (anonymous):

Can you please explain? I'm not sure how to figure it out

OpenStudy (anonymous):

for a function to be odd, f(-x)=-f(x) -f(x)=-(x^3-6x)=-x^3+6x f(-x)=(-x)^3-6(-x)=-x^3+6x f(-x)=-f(x), odd

OpenStudy (anonymous):

Oh, I see. Thank you

OpenStudy (unklerhaukus):

something like this